1. Amy borrowed $20,000 from her parents to open a bagel shop. She pays her parents a 5% yearly return on the money they lent her. Her other yearly fixed costs equal to $9,000. Her variable costs equal $30,000. In her first year, Amy sold 40,000 dozen at a price of $1.50 per dozen.

Amy's total costs equal

A. $40,000.

B. $39,000.

C. $59,000.

D. $50,000.
Answer #1

A. $40,000.

Return she pays on money = 5%*(20,000) = (0.05)*(20,000) = $1000
Total cost = Return she pays on money + Fixed cost + Variable cost = 1000 + 9000 + 30,000 = 40,000